Lisa is 9 years old. Marcus is 33 years old. How old will Lisa be when Marcus is three times as old as she is.
Answer by scott8148(6628)

M = L + 24 ___ M - 24 = L

M = 3L ___ L + 24 = 3L ___ 12 = L